OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (3, 6) and (1, −2)?

OpenStudy (anonymous):

two lines that are parallel have the same slope. the slope of the line passing through (3,6) and (1,-2) is m=(6+2)/(3-1)=8/2=4. so the required slope is also 4.